# SLAM&Render: A Benchmark for the Intersection Between Neural Rendering, Gaussian Splatting and SLAM https://samuel-cerezo.github.io/SLAM&Render

### v2

@@ Abstract
Models and methods originally developed for novel view synthesis and scene rendering, such as Neural Radiance Fields (NeRF) and Gaussian Splatting, are increasingly being adopted as representations in Simultaneous Localization and Mapping (SLAM). However, existing datasets fail to include the specific challenges of both fields, such as multi-modality and sequentiality in SLAM or generalization across viewpoints and illumination conditions in neural rendering. To bridge this gap, we introduce SLAM&Render, a novel dataset designed to benchmark methods in the intersection between SLAM and novel view rendering. It consists of 40 sequences with synchronized RGB, depth, IMU, robot kinematic data, and ground-truth pose streams. By releasing robot kinematic data, the dataset also enables the assessment of novel SLAM strategies when applied to robot manipulators. The dataset sequences span five different setups featuring consumer and industrial objects under four different lighting conditions, with separate training and test trajectories per scene, as well as object rearrangements. Our experimental results, obtained with several baselines from the literature, validate SLAM&Render as a relevant benchmark for this emerging research area.
I Introduction
Simultaneous Localization and Mapping, typically known by its acronym SLAM, is a key enabler for autonomous robot navigation and understanding. The goal is to simultaneously estimate the robot’s pose while creating a globally consistent representation or map of its surroundings, using only the data from the robot’s onboard sensors. After decades of intense research, several highly performing methods for geometric SLAM in rigid environments are available [1]. However, many research challenges remain [2]. SLAM differs from alternative approaches to 3D reconstruction (e.g., Structure from Motion [3]) by its online, real-time and multi-modal features, making it well-suited for robotic applications and requirements.
@@ I Introduction
Novel view synthesis, seemingly distant from robot SLAM, has experienced a resurgence due to groundbreaking research in neural radiance fields, also known as NeRFs [4]. Nevertheless, it has been quickly adopted as a map representation in visual [5, 6] and LiDAR [7] SLAM pipelines. On the other hand, Gaussian Splatting [8], developed as a much faster alternative to NeRFs, has had an even bigger impact in visual SLAM [9, 10, 11, 12]. Refer to the recent survey by Tosi et al. [13] for a comprehensive overview of references and in-depth discussions on how these two paradigms are influencing the field of SLAM.
Our work is motivated by this emerging trend intersecting visual SLAM, neural rendering and Gaussian splatting. Specifically, we argue that there is a lack of public datasets and benchmarks that explicitly address this area of research. Therefore, we introduce SLAM&Render, a novel dataset for benchmarking methods at the intersection of the aforementioned fields. It includes 40 sequences with synchronized RGB, depth, IMU, robot kinematic data, intrinsic and extrinsic camera parameters and ground-truth data. The sequences cover five setups with consumer and industrial objects captured under varying lighting conditions and include separate training/test trajectories and object rearrangements.
@@
All data is available online under the Creative Commons Attribution license (CC-BY 4.0) at the website. Additional information, example code and videos for visual inspection of the dataset is also included.
@@
II Related Work
@@
2014 DTU [14] NVS
2014 JIGSAWS [15] Motion Model.
2015 YCB object and Model set [16] Robotic Manip.
2016 EuRoC [17] SLAM
2017 Tanks and Temples [18] NVS
2018 TUM VI [19] Odom. (✓)
2019 OpenLORIS-Scene [20] SLAM
2019 Replica [21] SLAM & NVS
2021 Mip-NeRF360 [22] NVS
2023 ScanNet++ [23] SLAM & NVS
@@
TABLE I: Overview of commonly used datasets for SLAM and NVS, including available sensor modalities (RGB, Depth, IMU, Encoders), ground truth camera pose (GT), illumination variations (Ill.), scene rearrangements (Rearr.), and time synchronization (TS).
Publicly available datasets have played a crucial role in advancing research in SLAM and Novel View Synthesis (NVS). In Table I, we provide an overview of the most relevant datasets that are instrumental to these fields of study.
II-A Datasets for SLAM
@@
The EuRoC dataset [17] provides visual-inertial data from a Micro Aerial Vehicle, comprising two batches: one collected in an industrial environment for evaluating visual SLAM algorithms, and the other in a room for 3D environment reconstruction. Several challenges such as varying image brightness and dynamic scenes, are included. The TUM VI dataset [19] is designed for evaluating odometry algorithms. It includes RGB images, photometric calibration, and synchronized IMU data. Ground-truth poses at the start and end of the sequences are provided. The dataset also includes evaluations of state-of-the-art methods. Depth images are not present in this collection. OpenLORIS-Scene dataset [20] consists of visual-inertial and odometry data to benchmark SLAM and scene understanding algorithms in everyday scenarios. For data collection a wheeled robot equipped with a Realsense D435i camera, encoders and a LiDAR sensor was used. In the context of synthetic datasets, the Replica dataset [21] contains 18 photorealistic indoor reconstructions at room and building scales. It features dense 3D meshes, HDR textures, and reflective surfaces. However, neither changes in illumination or rearrangements of objects within the same scene are included.
II-B Datasets for Novel View Synthesis.
Despite the popularity of NeRFs, most datasets used for evaluation sample frames along a single camera trajectory, often lacking the diversity and complexity of real-world settings. DTU [14] was introduced for evaluating multi-view stereo techniques using a robotic setup with a structured light scanner. It contains a wide variety of scenes with different object types and lighting conditions. Tanks and Temples [18] was developed for image-based 3D reconstruction. It was captured in realistic, non-laboratory settings using an industrial laser scanner for ground-truth. It includes both indoor and outdoor scenes. Recently, Mip-NeRF360 [22] introduced a dataset for high-quality novel view synthesis. It captured camera trajectories around a central object in indoor and outdoor scenarios with detailed backgrounds. Then, ScanNet++ [23] was introduced to benchmark 3D scene understanding methods. It combines 3D laser scans, DSLR images, and RGB-D sequences. It includes semantic annotations for diverse and ambiguous scenarios and is the first to use independent camera trajectories for NVS evaluation. However, it does not contain IMU, kinematic data or controlled illumination changes.
In the context of datasets containing kinematic data, the YCB Object and Model Set [16] provides object manipulation data with joint poses but does not include visual data. On the other hand, the JIGSAWS dataset [15] includes kinematic data from the da Vinci surgical robot and stereoscopic images of surgical tasks, providing a rich data source to study human-robot interaction. However, the dataset purpose makes it very domain-specific.
Unlike the existing datasets, the one presented in this work offers a comprehensive and diverse set of sequences designed to benchmark methods in the intersection between SLAM and novel view rendering. The dataset includes 40 sequences with synchronized streams of RGB images, depth, IMU, robot kinematic data, and ground-truth poses. The sequences cover five distinct setups with both consumer and industrial objects, captured under four lighting conditions. Each scene includes separate trajectories for training and testing, and variations in object arrangement to evaluate generalization. To the best of our knowledge, this is the first dataset containing multi-sensor and kinematic data collected in sequences characterized by varying lighting conditions, object rearrangements, occlusions, and time-synchronization (see Table I).
III Notation
@@
III-A Convention for rigid transformations
Rigid transformations from a reference frame $A$ to another reference frame $B$ are defined as
$\bm{T}_{A}^{B}=\begin{bmatrix}\bm{R}_{A}^{B}&\bm{p}_{A}^{B}\\ \bm{0}_{1\times 3}&1\end{bmatrix}\in SE\left(3\right),$
where $\bm{R}_{A}^{B}\in SO\left(3\right)$ and $\bm{p}_{A}^{B}\in\mathbb{R}^{3}$ describe the orientation and the position of the frame $A$ with respect to the reference frame $B$, respectively. Such rigid transformation will transform a 3D point $\bm{p}^{A}\in\mathbb{R}^{3}$ ($\tilde{\bm{p}}^{A}$ in homogeneous coordinates) to the reference frame $B$ as $\tilde{\bm{p}}^{B}=\bm{T}_{A}^{B}\tilde{\bm{p}}^{A}$
III-B Forward kinematics
@@
$\bm{x}\left(t\right)=\bm{f}\left(\bm{q}\left(t\right)\right),$ (1)
where $\bm{f}\left(\cdot\right)$ is, in general, a nonlinear function allowing for the computation of the task variables starting from the robot configuration.
IV The SLAM&Render Dataset
The proposed dataset comprises 40 sequences that include both inertial and RGB-D data. These were recorded using an Intel RealSense D435i camera, mounted in an eye-to-hand configuration on a KUKA LBR iisy 3 R760 manipulator. Additionally, kinematic data were collected from the manipulator. In particular, the dataset contains the robot configurations $\bm{q}\left(t\right)$ along the motion and the corresponding flange poses computed using Eq. (1). An external Motion Capture System (MCS) was used to obtain the ground-truth camera poses. All sequences were recorded in an environment where the objects were placed on a black background, surrounded by black panels featuring pictures that display complex and colorful patterns (see Sec. IV-A). Moreover, the dataset contains train and test camera trajectories recorded under the following four different lighting conditions:
@@ IV The SLAM&Render Dataset
[…]
  • setup-1 and setup-2 include a set of common supermarket goods including transparent objects
  • setup-3 includes a different set of common supermarket goods including only opaque objects
  • setup-4 contains industrial objects
  • setup-5 contains industrial objects in a cluttered arrangement.
We additionally release a collection of images capturing a chessboard calibration pattern to calibrate the intrinsic and extrinsic parameters of the camera. The calibration procedure, together with the estimation of extrinsic and intrinsic parameters, is detailed in Sec. IV-E.
Our website offers a collection of tools for data association, retrieval from ROS2 bag files, alignment, and synchronization.
@@
IV-A Data acquisition
All RGB-D data were recorded at the optimal depth resolution [^2] https://dev.intelrealsense.com/docs/tuning-depth-cameras-for-best-performance (848×480) and full frame rate (30 Hz) of the Intel RealSense D435i camera. The IMU data were acquired at a rate of 210 Hz, capturing synchronized readings from the gyroscope and accelerometer. All the dataset sequences were collected through a system based on Ubuntu 22.04 and ROS2 Humble. We used the software library Intel RealSense SDK 2.0333https://www.intelrealsense.com/sdk-2/ and its ROS2 wrapper [^4] https://dev.intelrealsense.com/docs/ros2-wrapper to record the camera data. While the depth and color images are initially unaligned, as they are captured by two distinct sensors, we used the built-in feature of the Intel ROS2 wrapper to register the depth images to the color images. This is achieved by projecting the depth images to 3D and then back-projecting them into the color camera view. The joint angles of the robot manipulator were collected using ROS2 at a frequency of 25 Hz.
To record ground-truth data, we employed an OptiTrack Motion Capture System (MCS) composed of twelve PrimeX 13W cameras installed around the robot cell. These cameras track the position and orientation of passive markers at a frequency of 120 Hz. The markers are mounted on a circular structure between the camera and the robot flange (see Fig. 4). This structure was printed using an industrial-grade 3D printer, ensuring high accuracy while tracking the markers. Additionally, we have added two markers on the flange and one marker on the camera itself to enhance the asymmetry of the marker setup.
To minimize reflections from the surrounding environment, non-reflective black panels were installed around the perimeter of the robot cell. Pictures featuring complex and colorful patterns have been placed on the black panels not to oversimplify the novel view synthesis of NeRF-based algorithms. Further, the metal table supporting the robot manipulator was covered with a black cloth.
[…]
IV-B Dataset structure
[…]
.1 train. .2 rgb/. .3 1739371345.939813614.png. .3 1739371346.193981361.png. .3 …. .2 depth/. .3 1739371345.939813614.png. .3 1739371346.193981361.png. .3 …. .2 robot_data/. .3 joint_positions.txt. .3 flange_poses.txt. .2 associations.txt. .2 imu.txt. .2 groundtruth_raw.csv. .2 groundtruth.txt.

### v4

@@ Abstract
Models and methods originally developed for Novel View Synthesis, such as Neural Radiance Fields (NeRF) and Gaussian Splatting, are increasingly being adopted as representations in Simultaneous Localization and Mapping (SLAM). However, existing datasets fail to include specific research challenges of both fields, such as sequential processing and multi-modality in SLAM, or generalization across viewpoints and illumination changes in neural rendering. Additionally, datasets are often recorded using sensors that are handheld or mounted on drones or mobile robots, which complicates an accurate reproduction of sensor motions under different conditions. To bridge these gaps, we introduce SLAM&Render, a novel dataset designed to benchmark methods in the intersection between SLAM and Novel View Rendering. Recorded with a robot manipulator, it uniquely includes 40 sequences with time-synchronized RGB-D, IMU, robot kinematics and ground-truth pose streams. The dataset features five setups with consumer and industrial objects under four controlled lighting conditions, each with training and test trajectories with significant viewpoint changes. All sequences are static, with different levels of object rearrangements and occlusions. Our experimental results, obtained with several baselines from the literature, validate SLAM&Render as a relevant benchmark for this emerging research area. The dataset can be accessed through the following link: https://samuel-cerezo.github.io/SLAM&Render.
keywords:
Dataset, SLAM, Novel View Synthesis, Ground-truth
1 Introduction
[…]
@@ 1 Introduction
Novel View Synthesis (NVS), seemingly distant from robot SLAM, has experienced a resurgence due to groundbreaking research in Neural Radiance Fields (NeRFs) Mildenhall et al. (2021). NeRFs have been quickly adopted as a map representation in visual Sucar et al. (2021); Zhu et al. (2022) and LiDAR Deng et al. (2023) SLAM pipelines. On the other hand, Gaussian Splatting Kerbl et al. (2023), developed as a much faster alternative to NeRFs, has had an even bigger impact in visual SLAM Matsuki et al. (2024); Keetha et al. (2024). For a comprehensive overview of how these emerging paradigms are reshaping SLAM, see the recent survey by Tosi et al. Tosi et al. (2024).
Motivated by the growing convergence of visual SLAM, Neural Rendering, and Gaussian Splatting, and by the lack of benchmarks explicitly dedicated to it, we introduce SLAM&Render. Recorded with a robot manipulator, it ensures accurate reproductions of camera motions for different lighting conditions and scene setups, in order to study their effect separately. The sequences are divided into train and test, with significant viewpoint changes between them in order to assess generalization. The dataset provides 40 static sequences with synchronized RGB-D images, IMU readings, robot kinematic data, camera parameters, and ground-truth data. It comprises five setups with consumer and industrial objects under four controlled lighting conditions, showcasing object rearrangements with varying levels of occlusion.
@@
All data is available online under CC-BY 4.0 license at our website. Additional information, software tools, and videos for visual inspection of the data are also included.
@@
2 Related Work
Table 1: Overview of commonly used datasets for SLAM and NVS, including available sensor modalities (RGB, Depth, IMU, Encoders), ground-truth camera pose (GT), illumination variations (Ill.), scene rearrangements (Rearr.), and time synchronization (TS).
@@ 2 Related Work
2014 DTU Jensen et al. (2014) NVS
2014 JIGSAWS Gao et al. (2014) Motion Model.
2015 YCB object and Model set Calli et al. (2015) Robotic Manip.
2016 EuRoC Burri et al. (2016) SLAM
2017 Tanks and Temples Knapitsch et al. (2017) NVS
2018 TUM VI Schubert et al. (2018) Odom. (✓)
2019 OpenLORIS-Scene Shi et al. (2020) SLAM
2019 Replica Straub et al. (2019) SLAM & NVS
2020 Newer College Ramezani et al. (2020) SLAM & NVS
2021 Mip-NeRF360 Barron et al. (2022) NVS
2023 ScanNet++ Yeshwanth et al. (2023) SLAM & NVS
2025 Oxford Spires Tao et al. (2025) SLAM & NVS
@@
Publicly available datasets have played a crucial role in advancing research in SLAM and NVS. Table 1 provides an overview of the most relevant datasets in this field. Unless otherwise stated, the following datasets do not feature (i) robot kinematic data, (ii) full control of lighting conditions, (iii) reproducible sensor motions with high accuracy, and (iv) object rearrangements across sequences.
To our knowledge, SLAM&Render is the first dataset providing time-synchronized multi-sensor and kinematic data, object occlusions and rearrangements in a fully controlled recording setup (see Table 1).
2.1 Datasets for SLAM
@@
The EuRoC dataset Burri et al. (2016) provides visual–inertial data from a micro aerial vehicle organized into two batches: industrial hall sequences for SLAM evaluation, and indoor room sequences for 3D reconstruction. Several challenges, such as varying natural light conditions and dynamic scenes, are included. The TUM VI dataset Schubert et al. (2018) is designed for evaluating odometry algorithms. It contains RGB images, photometric calibration, synchronized IMU data, and ground-truth camera poses, while depth images are not included. The OpenLORIS-Scene dataset Shi et al. (2020) consists of visual–inertial, LiDAR and odometry data (i.e., encoder readings) from a mobile robot to benchmark SLAM and scene understanding in everyday scenarios. The Newer College Ramezani et al. (2020) and the Oxford Spires Tao et al. (2025) datasets provide visual-inertial, and LiDAR data from handheld devices. Both include time-synchronized data, calibration parameters for all the involved sensors, loop closures, and cluttered sequences observed from varying viewpoints. The Replica dataset Straub et al. (2019) includes 18 photorealistic indoor multi-scale reconstructions with dense meshes, rich textures, and reflective surfaces.
@@
2.2 Datasets for Novel View Synthesis
Despite the popularity of NeRFs, most datasets used for evaluation sample frames along a single camera trajectory, often lacking the diversity and complexity of real-world settings. DTU Jensen et al. (2014) was introduced to evaluate multi-view stereo techniques using a robotic setup for object 3D reconstruction tasks. It features systematic light control with a light scanner, however it addresses a very specific use case. In the same application field, Tanks and Temples Knapitsch et al. (2017) was captured in realistic indoor and outdoor settings using an industrial laser scanner for ground-truth collection. Recently, Mip-NeRF360 Barron et al. (2022) introduced a dataset for high-quality NVS. It captured camera trajectories around a central object in indoor and outdoor scenarios with detailed backgrounds. ScanNet++ Yeshwanth et al. (2023) benchmarks 3D scene understanding methods by combining 3D laser scans, high-quality images, and RGB-D streams with rich semantic annotations and independent camera trajectories for NVS.
Within datasets for robotics applications, the YCB Object and Model Set Calli et al. (2015) provides RGB-D images of several object manipulation sequences. In contrast, JIGSAWS Gao et al. (2014) dataset collects time-synchronized stereo images from an endoscopic camera, along with kinematic data from a surgical robot. However, it solely addresses applications in the field of surgery activities.
Unlike prior works, our dataset uniquely benchmarks SLAM, Neural Rendering, and Gaussian Splatting methods with 40 time-synchronized sequences featuring multimodal data, including RGB-D images, IMU readings, ground-truth, and robot kinematic data. Recorded with a robot manipulator it allows for accurately reproducing hand-eye camera motions, which are split in train and test trajectories. The five setups contain consumer and industrial objects under four controlled lighting conditions, which can generally have a negative impact on SLAM and Novel View Synthesis Ye et al. (2024); Martin-Brualla et al. (2021).
3 Notation
3.1 Convention for rigid transformations
A rigid transformation between two reference frames $A$ and $B$ is defined as $\mathbf{T}_{A}^{B}=\left[\mathbf{R}_{A}^{B},\,\mathbf{p}_{A}^{B};\,\mathbf{0}_{1\times 3},\,1\right]\in SE\left(3\right)$, where $\mathbf{R}_{A}^{B}\in SO\left(3\right)$ and $\mathbf{p}_{A}^{B}\in\mathbb{R}^{3}$ denote the rotation and translation of $A$ with respect to $B$, respectively. For any point $\mathbf{p}^{A}\in\mathbb{R}^{3}$, its coordinates in $B$ are $\tilde{\mathbf{p}}^{B}=\mathbf{T}_{A}^{B}\tilde{\mathbf{p}}^{A}$, where $\tilde{\mathbf{p}}^{A}$ and $\tilde{\mathbf{p}}^{B}$ are the homogeneous vectors of $\mathbf{p}^{A}$ and $\mathbf{p}^{B}$, respectively.
3.2 Forward kinematics
@@
$\bm{x}\left(t\right)={f}\left(\bm{q}\left(t\right)\right),$ (1)
where ${f}\left(\cdot\right)$ is, in general, a nonlinear function allowing for the computation of the task variables starting from the robot configuration.
4 The SLAM&Render Dataset
The proposed dataset comprises 40 static sequences that include both inertial and RGB-D data. These were recorded using an Intel RealSense D435i camera, mounted in an eye-to-hand configuration on a KUKA LBR iisy 3 R760 manipulator. The robot has been employed as a precise and reliable sensor carrier, ensuring consistent and reproducible camera trajectories. Additionally, kinematic data were collected from the manipulator. In particular, the dataset contains the robot configurations $\bm{q}\left(t\right)$ along the motion and the corresponding flange poses computed using (1). An external Motion Capture System (MCS) was used to obtain the ground-truth camera poses. All sequences were recorded in an environment where the objects were placed on a black background, surrounded by black panels featuring pictures that display complex and colorful patterns (see Sec. 4.1). Moreover, the dataset contains train and test camera trajectories recorded under the following four different lighting conditions:
@@ 4 The SLAM&Render Dataset
This enables our dataset to be used with classical SLAM approaches as well as newer Gaussian Splatting- and NeRF-based paradigms, which are particularly sensitive to lighting conditions and may require proper training to perform effectively. Therefore, this dataset feature ensures that SLAM algorithms are tested and trained in environments that reflect real-world variability, ultimately evaluating their performance and reliability in practical applications. Although the KUKA LBR iisy 3 R760 manipulator has a repeatability parameter [^1] The repeatability parameter gives a measure of the manipulator’s ability to return to a previously reached position. The smaller the value, the higher the repeatability. of ± $\pm\,0.1$ mm, we release the joint configurations and flange poses for all the sequences belonging to train and test trajectories. As a result, the camera pose estimation conducted by any SLAM algorithm on this dataset remains unaffected by potential variations in joint configurations. Furthermore, with smooth robot motions, the camera performs loop closures in both train and test trajectories (see Fig. 3). The sequences include scenes with both opaque and transparent objects, reflective and non-reflective materials, as well as varying occlusions resulting from the rearrangement of objects. The dataset features the following five distinct scenes (see Fig. 2):
  • setup-1 and setup-2: supermarket goods, including transparent objects;
  • setup-3: supermarket goods with only opaque objects;
  • setup-4: industrial objects;
  • setup-5: industrial objects in a cluttered arrangement.
We additionally release chessboard images for intrinsic and extrinsic calibration. The calibration procedure and resulting parameters are detailed in Sec. 4.5. Tools for data association, ROS2 bag data extraction, and synchronization can be found on our website.
@@
4.1 Data acquisition
All RGB-D data were recorded at the optimal depth [^2] https://dev.intelrealsense.com/docs/tuning-depth-cameras-for-best-performance (848×480) and full frame rate (30 Hz) of the Intel RealSense D435i camera. The IMU data were acquired at a rate of 210 Hz, capturing synchronized readings from the gyroscope and accelerometer. All the dataset sequences were collected through a system based on Ubuntu 22.04 and ROS2 Humble. We used the software library Intel RealSense SDK 2.0333https://www.intelrealsense.com/sdk-2/ and its ROS2 wrapper [^4] https://dev.intelrealsense.com/docs/ros2-wrapper to record the camera data. While the depth and color images are initially unaligned, as they are captured by two distinct sensors, we used the built-in feature of the Intel ROS2 wrapper to register the depth images to the color images. This is achieved by projecting the depth images to 3D and then back-projecting them into the color camera view. The joint angles of the robot manipulator were collected using ROS2 at a frequency of 25 Hz.
To record ground-truth data, we employed an OptiTrack MCS consisting of twelve PrimeX 13W cameras installed around the robot cell. These cameras track the position and orientation of passive markers at a frequency of 120 Hz. The markers are mounted on a circular structure between the camera and the robot flange (see Fig. 4). This structure was printed using an industrial-grade 3D printer, ensuring high accuracy while tracking the markers. Additionally, we have added two markers on the flange and one marker on the camera itself to enhance the asymmetry of the marker setup.
To minimize reflections from the surrounding environment, non-reflective black panels were installed around the perimeter of the robot cell. Pictures featuring complex and colorful patterns have been placed on the black panels to avoid oversimplifying the novel view synthesis of NeRF-based algorithms. Further, the metal table supporting the robot manipulator was covered with a black cloth. This recording setup was chosen to ensure high-quality, accurate ground-truth data, as environmental reflections and lighting variations within the robot cell degraded the ground-truth data in early recordings. The cold and warm lighting setups were obtained thanks to a Neewer 660 (50W) PRO LED lamp.
4.2 Dataset structure
Inspired by Sturm et al. (2012), our dataset is organized into sequences that share the same structure across all setups. For the sake of simplicity, consider the following setup-1 structure:
setup-1
    natural/
    warm/
    cold/
    dark/
    intrinsics.yaml
    extrinsics.yaml
Each lighting condition has a train and test sequence, which are organized equally. For example, the structure of the train sequence is reported below.
train
    rgb/
    1739371345.939813614.png
    1739371346.193981361.png
    depth/
    1739371345.939813614.png
    1739371346.193981361.png
    robot_data/
    joint_positions.txt
    flange_poses.txt
    associations.txt
    imu.txt
    groundtruth_raw.csv
    groundtruth.txt